a sporting goods store receives a shipment of 124 bags. two types of bags type a 38.50 type b 22.50 total cost is 3430 how many of each type?
Answer by stanbon(75887)

a sporting goods store receives a shipment of 124 bags.
two types of bags: type a 38.50 type b 22.50 total cost is 3430 how many of each type?
----
Quantity Eq.: a + b = 124
Value Eq.:::38.5a + 22.5b = 3430
-------------------------------------
Multiply thru 1st Eq by 385
Multiply thru 2nd Eq by 10
-------------------------------------
385a + 385b = 385*124
285a + 225b = 34300
-------------------------------
Subtract 2nd from 1st and solve for "b":
160b = 13440
b = 84 (# of type "b" bags
----
Since a + b = 124, a = 124-84 = 40 (# of type "a" bags)
=============================
Cheers,
Stan H.